OpenStudy (abbie):

8. which of the following equations is parallel to the line passing through A(-2, 7), B(-2,-2)

OpenStudy (abbie):

y=-2 y=x 3x=5 -2x=7y

OpenStudy (solomonzelman):

Ok, the lines that are paralyses to each other always have the same slope.

OpenStudy (solomonzelman):

parallel, I meant.

OpenStudy (abbie):

ok

OpenStudy (solomonzelman):

find the slope of the line that has points (-2, 7) (-2,-2) by plugging them into the slope formula, do you know the slope formula?

OpenStudy (abbie):

I got y=-2 but I feel like it's wrong?

OpenStudy (solomonzelman):

hold on, that can't be, b/c that has no slope, it is a vertical line.

OpenStudy (skullpatrol):

Draw a sketch.

OpenStudy (abbie):

ya m= y1-y2/x1-x2

OpenStudy (solomonzelman):

can you find the slope of the line that passes through (-2, 7) and (-2,-2) just plug these points into the slope formula.

OpenStudy (solomonzelman):

Understood?

OpenStudy (abbie):

ya wait though

OpenStudy (solomonzelman):

the slope is undefined, b/c it is a vertical line. sorry.

OpenStudy (solomonzelman):

if you plug the points into slope formula you will also see that.

OpenStudy (abbie):

that's what i thought. i was like the didnt ask me for the slope

OpenStudy (solomonzelman):

Yeah, that was my bad!

OpenStudy (solomonzelman):

there is no slope.

OpenStudy (skullpatrol):